A combined total of $20,000 is invested in two bonds that pay 2% and 8% simple interest. The annual interest is $1,180.00. How much is invested in each bond?
The amount invested at 2% is $_____
The amount invested at 8% is $_____

Answer by fractalier(6550)

Let us call what is invested at 2%, x.
The what is invested at 8% is 20000-x.
The setup looks like this then...
.02x+.08(20000-x)=1180
Now solve...
.02x+1600-.08x=1180
-.06x = -420
x = 7000 at 2%
20000-x = $13000 at 8%